# AnomalyGFM: Graph Foundation Model for Zero/Few-shot Anomaly Detection

### v1

@@ 3.4. Few-shot Prompt Tuning
To enable the few-shot prompt tuning in the scenarios where few-shot labeled normal nodes are available in the test/target graph, we utilize these limited nodes to have a prompt tunign of AnomalyGFM in the prototype learning module while keeping the GNN parameters frozen.
@@ 3.5. GAD Using AnomalyGFM
Inference on Very Large-scale Graphs. Most existing GAD methods typically require loading all nodes of the graph, which often leads to poor scalability on large graphs during inference. By learning the graph-agnostic prototypes, AnomalyGFM can generalize to very large-scale graphs through a subgraph-based inference approach. AnonalyGFM can effectively infer the anomaly score without considering the entire graph structure, eliminating the bottleneck of loading the full graph for GAD inference. This subgraph inference is also desired in privacy sensitive settings where we do not want to reveal the entire graph structure to the detection models. Specifically, AnomalyGFM operates in a way that given a test node $v_{i}$ and its subgraph $\mathcal{S}\left(v_{i}\right)$ with size $s$ are given, it first computes its residual feature ${\bf{r}}_{i}$ as:
@@
where ${\mathcal{S}}\left(v_{i}\right)$ is extracted using a random walk starting with the target node $v_{i}$. Then AnomalyGFM infer the anomaly score using the similarity of the residual feature to the normal and abnormal class prototypes as defined in Eq. (9).
@@ 4.1. Experimental Setup
Table 1. Key statistis of GAD datasets used in our experiments. Ano. indicate the number (rate) of anomalies in the graph and Sim. indicate the global average edge similarity.
@@
Facebook Social Networks 1,081 55,104 576 27(2.49%) 0.690
@@
Zero-shot Competing Methods. For the zero-shot setting, we evaluate AnomalyGFM against the state-of-the-art approaches from three main categories as follows (1) unsupervised methods: AnomalyDAE (Fan et al., 2020), CoLA (Liu et al., 2021b), TAM (Qiao and Pang, 2023a), and GADAM (Chen et al., 2024); (2) supervised methods: GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017), BWGNN (Tang et al., 2022), GHRN (Gao et al., 2023), and XGBGraph (Tang et al., 2023); and (3) general GFM methods, including GraphPrompt (Liu et al., 2023) for general graph tasks, and the one for zero-shot GAD UNPrompt (Niu et al., 2024). The detailed descriptions of each zero-shot competing method are provided in Appendix B.
Few-shot Competing Methods. In the few-shot setting, two generalist methods are included, i.e., GraphPrompt (Liu et al., 2023) and ARC (Liu et al., 2024). Detailed descriptions of each generalist competing method are provided in Appendix. B.
Evaluation Metrics. Following the previous studies (Liu et al., 2024; Tang et al., 2022; Qiao and Pang, 2023b), two widely-used metrics, AUROC and AUPRC, are used to evaluate the performance of all methods. For both metrics, a higher value denotes better performance. Moreover, for each method, we report the average performance for 3 independent runs with different random seeds. One-tailed Wilcoxon signed rank test is performed to evaluate the statistical significance of the performance of AnomalyGFM against each of its competing methods.
Implementation Details. For all competing methods and AnomalyGFM, the code is implemented with Pytorch 1.11 , PyG 2.1.0, DGL 1.0.1, and Python 3.8.19. All experiments are conducted on a Linux server with an Intel CPU (Intel Xeon Gold 6346 3.1GHz) and an Nvidia RTX3090 GPU with 24 GB GPU memory. The pre-train model is optimized using Adam optimizer (Kingma, 2014) with 300 epochs and a learning rate of $41e-$ 4 by default. The dimension of hidden space for representation and the prototype size are uniformly set to 300. The parameter $\alpha$ in the loss function is set to one by default. The setting of $\beta$ is determined based on the prior we gain from the global average similarity information provided in Table 1 (see Appendix C.2 for details). The $\mu$ and $\sigma$ of Gaussian distribution for both normal and abnormal class prototypes are set to zero and one by default. The size of the subgraph in subgraph inference is set to five by default. The implementation of competing methods is based on their publicly available official source code.
@@
AUROC 1-shot GraphPrompt (WebConf’23) 0.4216 0.4882 0.4223 0.2631 0.4811 0.5328 0.4086 0.6001 0.4000 0.4464 0.004
@@
5-shot GraphPrompt (WebConf’23) 0.4406 0.4900 0.6497 0.4726 0.5359 0.5381 0.4069 0.6012 0.4069 0.5046 0.003
@@
10-shot GraphPrompt (WebConf’23) 0.4321 0.4906 0.6314 0.7167 0.5367 0.5329 0.3826 0.6221 0.4260 0.5301 0.007
@@
AUPRC 1-shot GraphPrompt (WebConf’23) 0.0283 0.0680 0.0486 0.0426 0.1113 0.2321 0.0448 0.1108 0.0302 0.0796 0.012
@@
5-shot GraphPrompt (WebConf’23) 0.0285 0.0681 0.0892 0.0600 0.1661 0.2957 0.0327 0.1416 0.0360 0.1019 0.009
@@
10-shot GraphPrompt (WebConf’23) 0.0278 0.0681 0.0848 0.1427 0.1649 0.2922 0.0263 0.1421 0.0382 0.1096 0.007
@@ 4.2. Zero-shot GAD Performance
The AUROC and AUPRC results under the zero-shot setting are shown in Table 2 where the model is trained on Facebook and inference is made on other datasets without any further tuning. Similar results can be found when the models are trained on the Amazon dataset in Appendix C.1. From the table, we observe (1) AnomalyGFM consistently outperforms the three types of competing methods, including unsupervised methods, supervised methods, and generalist methods on most datasets. Specifically, AnomalyGFM achieves the best AUROC performance on six datasets and the best AUPRC performance on the five datasets, with up to 11% improvement in AUROC and 44% improvement in AUPRC when compared to the best-competing method. The main reason is that AnomalyGFM performs inference based on two discriminative graph-agonistic prototypes distilled from the node residual feature, supporting strong generalization across the datasets. (2) Unsupervised methods often fail to produce consistently good results across the datasets due to the fact that each of their model designs is devised in a way to focus on some specific abnormal patterns for the training graphs, making them struggle to generalize to the new graphs. (3) Most supervised methods perform well in zero-shot inference, particularly simple approaches like GCN, which achieves top performance on the Elliptic dataset. This can be attributed to their simple model design, avoiding heavy overfitting to the training graph. However, the performance still underperforms AnomalyGFM, demonstrating that the node representation residual is more generalized and our prototype alignment module provides a strong add-on to the binary loss function in our method. (4) Generalist methods, such as GraphPrompt, which are designed for general node/graph classification, struggle to perform well on GAD. This is due to the irregular nature of the abnormalities in different GAD datasets, hindering the applicability of the regular patterns/knowledge these generalist method learned from other tasks to GAD. AnomalyGFM outperforms UNPrompt on all datasets except Tolokers, which can be attributed to the node representation residual-based class prototypes offers more generalization power than the latent attribute predictability used in UNPrompt. Further, our statitical significance test results show that the superiority of AnomalyGFM over all competing methods is significant at the 90%/95% confidence level in both metrics.
@@ 4.3. Few-shot GAD Performance
To demonstrate the effectiveness in the few-shot scenarios, the comparison of AnomalyGFM to GraphPrompt (Liu et al., 2023) and ARC (Liu et al., 2024) is done under 1/5/10-shot. The AUROC and AUPRC results are shown in Table 3. In the few-shot GAD scenario, AnomalyGFM outperforms these two generalist methods on most datasets in both AUROC and AUPRC. GraphPrompt continues to underperform AnomalyGFM since the presence of only one class of samples renders its fine-tuning ineffective, leading to suboptimal performance. ARC also leverages residual features and is specifically designed for few-shot GAD datasets achieving the best performance on the Question and the second performance on most datasets. Compared to ARC, AnomalyGFM achieves better sample-efficiency tuning due to the learning of class-level prototypes that are agnostic to different domains of graphs, enhancing the generalization across the graph, thereby yielding the best performance on the eight datasets.
4.4. Scale up to Very Large Graphs
We choose two large-scale graphs T-Finance (with a large number of edges) and T-Social (with a large number of nodes)
@@ 4.4. Scale up to Very Large Graphs
to evaluate the generalization of AnomalyGFM on very large graphs. The results are shown in Table 4. We can observe the AnomalyGFM outperforms the unsupervised methods and supervised methods on these datasets. This is mainly because the learned data-independent prototypes can effectively measure the abnormality of node representation residual, even in the presence of subgraphs. The deviations of the target node from the contextual node are still preserved in the subgraph, thereby enabling effective subgraph-bsed inference.
Table 5. Abalation study results for AnomalyGFM and its variants.
@@ 4.5. Ablation Study
To verify the effectiveness of each component of AnomalyGFM, we designed three variants including (i) BCE which replaces the similarity measurement with the predicted probability from the BCE loss function for anomaly scoring. (ii) BCE Residual (BCE-R) which firstly removes the alignment loss and applies the BCE loss on the residual features to differentiate normal and abnormal nodes. Then, the predicted probability is used as the anomaly score. (iii) Feature Alignment (FA) which replaces the alignment on the residual feature with the alignment on the original feature. The results are shown in Table 5. FA achieves the highest AUROC on the Elliptic and T-Finance datasets, suggesting that learning prompts from the original features can also be effective as the original feature also maintains anomaly discriminability but they aren’t in the unified feature space leading to sub-optimal performance. BCE does not perform well except on the small dataset Disney. It directly uses the predicted probability as the anomaly score, which makes it struggle to generalize to new graphs due to the differences between the training and test sets. BCE-R does not achieve good performance compared to BCE. The reason is that although the residual features are discriminative, many samples located on the decision boundaries after being mapped to a unified feature space may not be effectively scored by the supervised classifier. In contrast, aligning the learnable prototype with the average node representation residual and using similarity measurement for scoring can effectively discriminate the normal and abnormal nodes.
@@ 4.6. Hyperparameter Sensitivity Analysis
We evaluate the sensitivity of AnomalyGFM w.r.t the size of the prototype $T$, hyperparameter $\alpha$, and commonality dimension $d^{\prime}$. Additional sensitivity analyses are in Appendix C.2.
@@
Sensitivity of AnomalyGFM w.r.t common dimensionality $d^{\prime}$. To further demonstrate AnomalyGFM’s generalization, we evaluate AnomalyGFM under different $d^{\prime}$. The results are shown in Fig. 6. AnomalyGFM outperforms all the competing methods under different dimensions. The main reason is that the deviation between connected nodes is still preserved with different $d^{\prime}$ which can be efficiently measured by the graph-agnostic prototypes. The AUPRC results are provided in Appendix C.2.
Figure 6. AUROC of AnomalyGFM w.r.t $d^{\prime}$

### v2

@@ 3.4. Few-shot Prompt Tuning
To enable the few-shot prompt tuning in the scenarios where few-shot labeled normal nodes are available in the test/target graph, we utilize these limited nodes to have a prompt tuning of AnomalyGFM in the prototype learning module while keeping the GNN parameters frozen.
@@ 3.5. GAD Using AnomalyGFM
Inference on Very Large-scale Graphs. Most existing GAD methods typically require loading all nodes of the graph, which often leads to poor scalability on large graphs during inference. By learning the graph-agnostic prototypes, AnomalyGFM can generalize to very large-scale graphs through a subgraph-based inference approach. AnonalyGFM can effectively infer the anomaly score without considering the entire graph structure, eliminating the bottleneck of loading the full graph for GAD inference. This subgraph inference is also desired in privacy-sensitive settings where we do not want to reveal the entire graph structure to the detection models. Specifically, AnomalyGFM operates in a way that given a test node $v_{i}$ and its subgraph $\mathcal{S}\left(v_{i}\right)$ with size $s$ are given, it first computes its residual feature ${\bf{r}}_{i}$ as:
@@
where ${\mathcal{S}}\left(v_{i}\right)$ is extracted using a random walk starting with the target node $v_{i}$. Then, AnomalyGFM infers the anomaly score using the similarity of the residual feature to the normal and abnormal class prototypes as defined in Eq. (9).
@@ 4.1. Experimental Setup
Table 1. Key statistics of GAD datasets used in our experiments. Ano. indicates the number (rate) of anomalies in the graph and Sim. indicates the global average edge similarity.
@@
Facebook Social Networks 1,081 55,104 576 25(2.31%) 0.690
@@
Zero-shot Competing Methods. For the zero-shot setting, we evaluate AnomalyGFM against the state-of-the-art approaches from three main categories as follows (1) unsupervised methods: AnomalyDAE (Fan et al., 2020), CoLA (Liu et al., 2021b), TAM (Qiao and Pang, 2023), and GADAM (Chen et al., 2024); (2) supervised methods: GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017), BWGNN (Tang et al., 2022), GHRN (Gao et al., 2023), and XGBGraph (Tang et al., 2023); and (3) general GFM methods, including GraphPrompt (Liu et al., 2023) for general graph tasks, and zero-shot GAD UNPrompt (Niu et al., 2024).
Few-shot Competing Methods. In few-shot setting, three generalist methods are icluded, i.e.GPPT (Sun et al., 2022), GraphPrompt (Liu et al., 2023) and ARC (Liu et al., 2024).
Evaluation Metrics. Following the previous studies (Liu et al., 2024; Tang et al., 2022; Qiao and Pang, 2023), two widely-used metrics, AUROC and AUPRC, are used to evaluate the performance of all methods. For both metrics, a higher value denotes better performance. Moreover, for each method, we report the average performance for 3 independent runs with different random seeds. One-tailed Wilcoxon signed rank test is performed to evaluate the statistical significance of the performance.
Implementation Details. For all competing methods and AnomalyGFM, the code is implemented with Pytorch 1.11 , PyG 2.1.0, DGL 1.0.1, and Python 3.8.19. All experiments are conducted on a Linux server with an Intel CPU (Intel Xeon Gold 6346 3.1GHz) and an Nvidia RTX3090 GPU with 24 GB GPU memory. The pre-train model is optimized using the Adam optimizer (Kingma, 2014) with 300 epochs and a learning rate of $41e-$ 4 by default. The dimension of hidden space for representation and the prototype size are uniformly set to 300. The parameter $\alpha$ in the loss function is set to one by default. The setting of $\beta$ is determined based on the prior we gain from the global average similarity information provided in Table 1 (see Sec. 6 for details). The $\mu$ and $\sigma$ of the Gaussian distribution for both normal and abnormal class prototypes are set to zero and one by default. The size of the subgraph in subgraph inference is set to five by default. The implementation of competing methods is based on their publicly available official source code.
@@
AUROC 1-shot GPPT (KDD’22) 0.5000 0.5303 0.4997 0.5010 0.5000 0.5061 0.4921 0.6162 0.3647 0.5011 0.003
GraphPrompt (WebConf’23) 0.4216 0.4882 0.4223 0.2631 0.4811 0.5328 0.4086 0.6001 0.4000 0.4464 0.004
@@
5-shot GPPT (KDD’22) 0.5000 0.5098 0.5000 0.5051 0.5000 0.5181 0.4959 0.5736 0.2609 0.4848 0.003
GraphPrompt (WebConf’23) 0.4406 0.4900 0.6497 0.4726 0.5359 0.5381 0.4069 0.6012 0.4069 0.5046 0.003
@@
10-shot GPPT (KDD’22) 0.5000 0.5087 0.4769 0.5023 0.5000 0.4971 0.5047 0.4212 0.5539 0.4961 0.003
GraphPrompt (WebConf’23) 0.4321 0.4906 0.6314 0.7167 0.5367 0.5329 0.3826 0.6221 0.4260 0.5301 0.007
@@
AUPRC 1-shot GPPT (KDD’22) 0.0333 0.0766 0.0488 0.0687 0.1453 0.2204 0.0294 0.1239 0.0432 0.0877 0.003
GraphPrompt (WebConf’23) 0.0283 0.0680 0.0486 0.0426 0.1113 0.2321 0.0448 0.1108 0.0302 0.0796 0.012
@@
5-shot GPPT (KDD’22) 0.0333 0.0692 0.0504 0.0716 0.1453 0.2265 0.0297 0.1127 0.0365 0.0861 0.004
GraphPrompt (WebConf’23) 0.0285 0.0681 0.0892 0.0600 0.1661 0.2957 0.0327 0.1416 0.0360 0.1019 0.009
@@
10-shot GPPT (KDD’22) 0.0334 0.0691 0.0526 0.0698 0.1453 0.2178 0.0301 0.0905 0.0511 0.0844 0.004
GraphPrompt (WebConf’23) 0.0278 0.0681 0.0848 0.1427 0.1649 0.2922 0.0263 0.1421 0.0382 0.1096 0.007
@@ 4.2. Zero-shot GAD Performance
The AUROC and AUPRC results under the zero-shot setting are shown in Table 2 where the model is trained on Facebook and inference is made on other datasets without any further tuning. Similar results can be found when the models are trained on the Amazon dataset in Appendix B.2. From the table, we observe (1) AnomalyGFM consistently outperforms the three types of competing methods, including unsupervised methods, supervised methods, and generalist methods on most datasets. Specifically, AnomalyGFM achieves the best AUROC performance on six datasets and the best AUPRC performance on the five datasets, with up to 11% improvement in AUROC and 44% improvement in AUPRC when compared to the best-competing method. The main reason is that AnomalyGFM performs inference based on two discriminative graph-agonistic prototypes distilled from the node residual feature, supporting strong generalization across the datasets. (2) Unsupervised methods often fail to produce consistently good results across the datasets due to the fact that each of their model designs is devised in a way to focus on some specific abnormal patterns for the training graphs, making them struggle to generalize to the new graphs. (3) Most supervised methods perform well in zero-shot inference, particularly simple approaches like GCN, which achieves top performance on the Elliptic dataset. This can be attributed to their simple model design, avoiding heavy overfitting to the training graph. However, the performance still underperforms AnomalyGFM, demonstrating that the node representation residual is more generalized and our prototype alignment module provides a strong add-on to the binary loss function in our method. (4) Generalist methods, such as GraphPrompt, which are designed for general node/graph classification, struggle to perform well on GAD. This is due to the irregular nature of the abnormalities in different GAD datasets, hindering the applicability of the regular patterns/knowledge these generalist method learned from other tasks to GAD. AnomalyGFM outperforms UNPrompt on all datasets except Tolokers, which can be attributed to the node representation residual-based class prototypes offers more generalization power than the latent attribute predictability used in UNPrompt. Further, our statistical significance test results show that the superiority of AnomalyGFM over all competing methods is significant at the 90%/95% confidence level in both metrics.
@@ 4.3. Few-shot GAD Performance
To demonstrate the effectiveness in the few-shot scenarios, the comparison of AnomalyGFM to GraphPrompt (Liu et al., 2023) and ARC (Liu et al., 2024) is done under 1/5/10-shot. The AUROC and AUPRC results are shown in Table 3. In the few-shot GAD scenario, AnomalyGFM outperforms these two generalist methods on most datasets in both AUROC and AUPRC. GraphPrompt continues to underperform AnomalyGFM since the presence of only one class of samples renders its fine-tuning ineffective, leading to suboptimal performance. ARC also leverages residual features
@@
and is specifically designed for few-shot GAD datasets achieving the best performance on the Question and the second performance on most datasets. Compared to ARC, AnomalyGFM achieves better sample-efficiency tuning due to the learning of class-level prototypes that are agnostic to different domains of graphs, enhancing the generalization across the graph, thereby yielding the best performance on the eight datasets. Tuning the abnormal prototype is also possible when a few labeled anomalies are available and the results are provided in App. B.1.
4.4. Scale up to Very Large Graphs
We choose two large-scale graphs T-Finance (with a large number of edges) and T-Social (with a large number of nodes) to evaluate the generalization of AnomalyGFM on very large graphs. The results are shown in Table 4. We can observe that AnomalyGFM outperforms the unsupervised methods and supervised methods on these datasets. This is mainly because the learned data-independent prototypes can effectively measure the abnormality of node representation residual, even in the presence of subgraphs. The deviations of the target node from the contextual node are still preserved in the subgraph, thereby enabling effective subgraph-based inference.
Table 5. Ablation study results for AnomalyGFM and its variants.
@@ 4.5. Ablation Study
To verify the effectiveness of each component of AnomalyGFM, we designed three variants including (i) BCE which replaces the similarity measurement with the predicted probability from the BCE loss function for anomaly scoring. (ii) BCE Residual (BCE-R) which firstly removes the alignment loss and applies the BCE loss on the residual features to differentiate normal and abnormal nodes. Then, the predicted probability is used as the anomaly score. (iii) Feature Alignment (FA) which replaces the alignment on the residual feature with the alignment on the original feature. The results are shown in Table 5. FA achieves the highest AUROC on the Elliptic and T-Finance datasets, suggesting that learning prompts from the original features can also be effective as the original feature also maintains anomaly discriminability but they aren’t in the unified feature space leading to sub-optimal performance. BCE does not perform well except on the small dataset Disney. It directly uses the predicted probability as the anomaly score, which makes it struggle to generalize to new graphs due to the differences between the training and test sets. BCE-R does not achieve good performance compared to BCE. The reason is that although the residual features are discriminative, many samples located on the decision boundaries after being mapped to a unified feature space may not be effectively scored by the supervised classifier. In contrast, aligning the learnable prototype with the average node representation residual and using a similarity measurement for scoring can effectively discriminate the normal and abnormal nodes.
@@ 4.6. Hyperparameter Sensitivity Analysis
We evaluate the sensitivity of AnomalyGFM w.r.t the size of the prototype $T$, hyperparameter $\alpha$, and commonality dimension $d^{\prime}$. Additional sensitivity analyses are in Appendix B.3.
@@
Sensitivity of AnomalyGFM w.r.t common dimensionality $d^{\prime}$. To further demonstrate AnomalyGFM’s generalization, we evaluate AnomalyGFM under different $d^{\prime}$. The results are shown in Fig. 6. AnomalyGFM outperforms all the competing methods under different dimensions. The main reason is that the deviation between connected nodes is still preserved with different $d^{\prime}$ which can be efficiently measured by the graph-agnostic prototypes.
Figure 6. AUROC and AUPRC of AnomalyGFM w.r.t $d^{\prime}$
Impact of $\beta$ in anomaly scoring. The performance under different $\beta$ is shown in Fig. 7. As shown in Table 9 in App. B.3, the set of $\beta$ values are determined based on the prior we gain from the global average similarity. Specifically, the hyperparameter in anomaly scoring $\beta$ is set to zero by default for graphs with high global similarity such as Amazon and Yelp, and four for those with low global similarity such as Ellitipic and T-Finance, see the detailed global similarity information in Table 1. In the few-shot scoring, $\beta$ is set to $0.5$ by default for graphs with high global similarity. The main reason for these settings is that for graphs with high global average similarity, where node features are relatively similar, the deviations in local neighborhood would be small, and the fluctuations in scoring based on the normal prototype may affect the overall anomaly score. Therefore, we solely use the abnormal class prototype or assign a smaller weight for the normal class prototype in scoring. For nodes with less similar features, we use a mix of normal and abnormal prototypes for scoring. In the few-shot setting, the labeled normal nodes help refine the normal prototypes, allowing the anomaly scores to be computed using both normal and abnormal prototypes.
Figure 7. AUROC and AUPRC of AnomalyGFM w.r.t $\beta$.
Table 6. Training and inference time (seconds) comparison.
Methods AnomalyDAE TAM GAT BWGNN UNPrompt AnomalyGFM
Training Time 86.04 479.70 2.43 4.86 2.08 7.45
Inference Time 170.46 252.52 91.88 98.33 19.47 11.00
5. Time Complexity Analysis
In this section, we analyze the time complexity of AnomalyGFM. We build the GCN as the backbone of AnomalyGFM which takes $O(mdh)$, where $m$ is the number of non-zero elements in matrix $\bf A$ (the number of edge in graph), $d$ is the dimension of representation, and $h$ is the number of feature maps. The computation of the residual feature takes $O(N^{2}d)$. The MLP layer mapping the representation to the anomaly score takes $O(Nd)$. The feature mapping of normal class prototype and abnormal class prototype takes $O(2d^{2})$. The alignment loss between residual features and prototypes takes $O(2Nd)$. Thus, the overall complexity of the pre-training of AnomalyGFM is $O(mdh+N^{2}d+3Nd+2d^{2})$. In the inference, we employ the residual feature’s similarity with two prototypes which takes $O(2Nd)$. Therefore, the overall complexity of inference of AnomalyGFM is $O(mdh+N^{2}d+2d^{2}+2Nd)$. The complexity of inference on large-scale graph takes $O\left({K\left({mdh+{n^{2}}d+2{d^{2}}+2nd}\right)}\right)$ with $n$ as subgraph size and $K$ as the number of subgraphs. It is significantly lower than other methods, as they require loading entire graphs and performing reconstruction or affinity calculations, leading to higher space/time complexity. In Table 6, we report the running time comparison on Facebook and inference time on YelpChi-all. AnomalyGFM requires much less training and inference time compared to TAM and AnomalyDAE, demonstrating the efficiency. For inference, AnomalyGFM is the most efficient method because scoring relies solely on measuring the similarity between node representation residuals and prototypes.